A car started out from Memphis toward Little Rock at the rate of 60km/h. A second car left from the same point 2 hours later and drove along the same route at 75km/h. How long did it take the second car to catch the first car. Use the formula d= RxT, and also make a chart.
Answer by checkley79(3341)

D=RT
D=60T FOR THE FIRST CAR.
D=75(T-2) FOR THE SECOND CAR.
60T=75(T-150)
60T=75T-150
60T-75T=-150
-15T=-150
T=-150/-15
T=10 HOURS AFTER THE FIRST CAR LEAVES THEY WILL MEET.
PROOF:
60*10=75(10-2)
600=75*8
600=600
10-2=8 HOURS FOR THE SECOND CAR TO CATCH THE FIRST CAR.